OpenStudy (zmudz):

How many solutions are there to the equation \(\left| x^2 - 6x \right| = 9\)? I put 2 but its not right.

OpenStudy (campbell_st):

well you need to solve \[x^2 - 6x = 9~~~and~~~~x^2 - 6x = -9\] then check to see if the solutions make sense by substituting

OpenStudy (lynfran):

|x^2-6x|=9 x^2-6x=+ 9......, x^2-6x=-9 x^2-6x-9=0 and x^2-6x+9=0 now use the quadratic formula to find solution or if it can be factor...